# Model Already Knows the Best Noise: Bayesian Active Noise Selection via Attention in Video Diffusion Model

### v1

@@ 3.3 Layer Selection via Cumulative BANSA Correlation
${\widehat{\text{BANSA-E}}}_{\leq d}(\mathbf{z}_{i},c,t):=\frac{1}{d}\sum_{l=1}^{d}{\text{BANSA-E}}^{(l)}(\mathbf{z}_{i},c,t).\vspace{-0.5em}$ (9)
To determine $d^{*}$, we compute the Pearson correlation (46) between ${\widehat{\text{BANSA-E}}}_{\leq d}$ and the full-layer average ${\widehat{\text{BANSA-E}}}_{\leq L}$, and select the smallest $d$ satisfying:
$\text{Corr}\left({\widehat{\text{BANSA-E}}}_{\leq d},{\widehat{\text{BANSA-E}}}_{\leq L}\right)\geq\tau,$ (10)
with $\tau=0.7$ in our experiments. We validate this procedure using 100 prompts and 10 noise seeds across CogVideoX-2B, and CogVideoX-5B. As shown in Figure 5, the correlation stabilizes at layer 14 in CogVideoX-2B, and 19 in CogVideoX-5B. We therefore set $d^{*}$ accordingly and define the BANSA score as ${\widehat{\text{BANSA-E}}}_{\leq d^{*}}$ to guide noise selection, as summarized in Algorithm 1.
This layer selection procedure provides a lightweight and model-specific approximation of the full BANSA score. Since $d^{*}$ can be predefined for each model, it does not interfere with the noise sampling process and introduces no runtime cost during generation. As shown in the Appendix, ${\widehat{\text{BANSA-E}}}_{\leq d^{*}}$ closely approximates the full-layer score and achieves comparable generation quality across all models.
@@ 4 Experiments
Experimental Setting. We evaluate ANSE on two representative text-to-video (T2V) diffusion models: CogVideoX-2B and CogVideoX-5B (9), chosen for their strong spatiotemporal modeling capabilities grounded in real-world dynamics. This setup enables rigorous evaluation of noise selection where attention and coherence are critical. We follow the official DDIM sampling protocol with 50 denoising steps for both models. Quantitative results for noise prior based approaches such as Freeinit (1) and FreqPrior (2) are omitted, as they are not officially supported on CogVideoX and incur 3$\times$ inference cost. Nonetheless, ANSE is orthogonal and can be combined with these methods for further gains. We use a noise pool of size $M{=}10$ with $K{=}10$ stochastic forward passes per noise, and apply Bernoulli-masked attention with a masking probability $p{=}0.2$. Additional details are in the Appendix. All experiments are run on NVIDIA H100 GPUs.
Evaluation Metric. To evaluate the impact of ANSE, we use VBench (47), a perceptually grounded benchmark for text-to-video generation. VBench reports two high-level metrics—quality score and semantic score—which are combined into a total score via weighted averaging and normalized to a 0–100 scale. Each score is a composite metric: the quality score is derived from 7 perceptual dimensions including subject consistency, background consistency, temporal flickering, motion smoothness, dynamic degree, aesthetic quality, and imaging quality; the semantic score comprises 9 alignment-related criteria such as object class, multiple objects, human action, color, spatial relationship, scene, temporal style, appearance style, and overall consistency. This decomposition ensures that VBench comprehensively assesses both visual fidelity and semantic alignment. For each configuration (with and without BANSA), we generate 4,730 videos to ensure statistical reliability.
Table 1: Quantitative results on VBench using CogVideoX-2B and -5B. ANSE consistently improves quality, semantic alignment, and total score. Scores for noise prior methods (1; 2) are omitted as they are not officially supported on CogVideoX and require costly re-implementation (3$\times$ inference), making large-scale evaluation impractical.
@@
Model Backbone Method Quality Score Semantic Score Total Score Inference Time
CogVideoX-2B (9) Vanilla 82.08 76.83 81.03 247.8
+ Ours 82.56(+0.48) 78.06(+1.23) 81.66(+0.63) 269.3(+8.68%)
CogVideoX-5B (9) Vanilla 82.53 77.50 81.52 1223.5
+ Ours 82.70(+0.17) 78.10(+0.60) 81.71(+0.25) 1392.1(+13.78%)
@@
Figure 4: Qualitative comparison of CogVideoX variants with and without ANSE. Results from CogVideoX-2B are shown in the first two rows; the rest show results from CogVideoX-5B. With ANSE, videos exhibit improved visual quality, better text alignment, and smoother motion transitions compared to the baseline.
Quantitative Comparison. As shown in Table 1, ANSE consistently improves performance across both CogVideoX models. On CogVideoX-2B, the total VBench score increases from 81.03 (Vanilla) to 81.66 with ANSE, driven by gains in quality (+0.48) and semantic alignment (+1.23). On the larger CogVideoX-5B, ANSE also improves all metrics: quality increases from 82.53 to 82.70 (+0.17), semantic score from 77.50 to 78.10 (+0.60), and total score from 81.52 to 81.72 (+0.25). These results demonstrate that ANSE effectively enhances both perceptual fidelity and semantic alignment, even on large-scale, temporally grounded video diffusion models. The consistent gains across both model scales highlight the robustness and generalizability of ANSE in selecting high-quality noise seeds under varying architectural complexities.
Qualitative Comparison. Figure 4 shows qualitative results on CogVideoX-2B and -5B with and without ANSE. Our method improves semantic fidelity, motion portrayal, and visual clarity across diverse prompts. For example, in "exploding" and ANSE captures key semantic transitions—generating visible explosions and preserving temporal continuity. In "vacuuming", the subject remains static in the baseline but exhibits purposeful motion with ANSE.
On CogVideoX-5B, similar improvements are evident. In "koala playing the piano" and "zebra running", ANSE generates anatomically coherent bodies with expressive motion. These results demonstrate ANSE’s ability to enhance spatial-temporal fidelity and generalize to high-capacity video diffusion models.
Computational Cost. As shown in Table 1, ANSE increases inference time by only +8% on CogVideoX-2B and +13% on CogVideoX-5B, measured in denoising steps. This overhead stems from noise seed evaluation but does not affect the sampling process, so memory usage remains unchanged. In contrast, prior methods such as Freeinit and FreqPrior require three full sampling passes, resulting in a 200% increase in inference time. While these methods have not been officially implemented, making direct comparisons challenging, ANSE reduces inference cost by approximately 64% while achieving comparable or superior generation quality.
5 Ablation Study and Analysis
@@
Comparison of Acquisition Functions. We compare BANSA with alternative acquisition strategies for noise seed selection using the CogVideoX-2B model. As shown in Table 3, we evaluate random sampling, entropy-based selection, and two BANSA variants: BANSA (B), which uses Bernoulli masking, and BANSA (D), which introduces Dropout-based stochasticity. While all methods improve over the baseline, BANSA (B) consistently achieves the highest scores across quality, semantic, and total metrics. This confirms that injecting uncertainty through Bernoulli masking is more effective than dropout, which is commonly used in Bayesian acquisition functions such as BALD. The result highlights the importance of modeling attention-level uncertainty in a manner that reflects the structure of the underlying model.
Effect of Ensemble Size $K$. We assess the impact of the number of stochastic forward passes $K$ on subject and background consistency, again using CogVideoX-2B. As shown in Table 3, both metrics improve steadily as $K$ increases from 1 to 10, suggesting that larger ensembles yield more robust and stable noise evaluations. Performance saturates at $K=10$, which we adopt as the default throughout all experiments.
Effect of Noise Pool Size $M$. We analyze the role of noise pool size $M$, which determines the diversity of candidate seeds assessed by BANSA. While a larger $M$ increases the likelihood of discovering high-quality seeds, it also raises inference cost. As shown in Figure 6, performance saturates around $M=10$ for CogVideoX-2B and -5B. We set these values as defaults for each model.
@@ 6 Discussion and Limitations
Our method focuses on noise seed selection through model uncertainty estimation, yet it has notable limitations. As shown in Figure 7, even seeds with low BANSA scores, which indicate high model confidence, can produce unnatural generations. This suggests that while ANSE effectively identifies promising initial seeds, it does not directly affect the generation process itself. Moreover, there remains a gap between estimated uncertainty and perceptual quality. Although BANSA reliably captures attention-level uncertainty, it may not fully reflect semantic or aesthetic aspects. While generating multiple candidates per seed and selecting based on strong quality metrics would be ideal, this approach is computationally prohibitive. We thus view BANSA as a practical surrogate for such strategies. Future work could further enhance performance by integrating it with information-theoretic refinement or active learning methods.
@@ 7 Conclusion
We present ANSE, a framework for active noise selection in video diffusion models. It is built around BANSA, an acquisition function that uses attention-derived uncertainty to identify noise seeds that promote confident and consistent attention, which are indicative of high-quality generation. BANSA adapts the BALD principle to the generative setting by operating in the attention space. To enable efficient deployment, we introduce a stochastic approximation using Bernoulli-masked attention and a lightweight layer selection method. Experiments across multiple T2V backbones show that ANSE improves video quality and prompt alignment with little to no increase in inference time. This work highlights the potential of inference-time noise selection guided by internal model signals. By extending active learning concepts beyond training-time data selection, ANSE enables model-aware decisions during inference and enhances generation without retraining. Our approach introduces a new inference-time scaling paradigm, where performance is improved not by modifying the model or increasing sampling steps, but through informed seed selection.
@@ Appendix A Supplementary Section
  • Implementation details of the BANSA score in Section B.
  • Proof of Proposition 1 from the main paper regarding the BANSA Zero condition in Section C.
  • Implementation details of ANSE and evaluation metrics in Section D.
  • Further explanation of layer selection through cumulative BANSA correlation in Section E.
  • Additional ablation studies including full-layer BANSA Score analysis in Section F and temporal scope effects in Section G.
  • Additional qualitative results demonstrating the impact of BANSA Score in Section H.
Appendix B Implementation Details of BANSA Score
Score Definition vs. Implementation.
In the main paper, we define the BANSA score as the difference between the mean of entropies and the entropy of the mean:
$\text{BANSA}(\mathbf{z},c,t):=\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)})-\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\right),$
This corresponds to the negative of the mutual information formulation used in BALD (29). But, for practical and interpretability reasons, we adopt the BALD-style computation in our implementation:
$\text{BANSA}(\mathbf{z},c,t):=\mathcal{H}\left(\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\right)-\frac{1}{K}\sum_{k=1}^{K}\mathcal{H}(A^{(k)}),$
which yields a non-negative score. This implementation allows easier visualization and interpretation, where larger values correspond to higher disagreement (i.e., uncertainty) across stochastic attention maps. We adopt this sign convention to align with standard BALD implementations, where mutual information is expressed as a non-negative measure of epistemic uncertainty. All figures and tables in the paper (e.g., correlation plots and performance curves) reflect this convention, showing positive-valued BANSA scores.
Semantic consistency.
Although the mathematical sign differs between the definition and implementation, the semantic meaning and selection behavior are strictly preserved. In both cases, we minimize the BANSA score to select noise seeds with low attention uncertainty. The relative ordering of scores remains unchanged, and the interpretation of the score as a measure of model disagreement is fully retained.
For clarity, we note that the proof of Proposition 1 in the following section is also presented under this BALD-style convention.
Appendix C Proof of Proposition 1

### v2

@@ 4 Analysis of BANSA-Selected Noise
[…]
To demonstrate this prompt-dependent behavior, We select a high-BANSA (worst-performing) seed from prompt A and applied it to prompts B and C (Table 1) and measures Subject and Background Consistency. The seed degraded quality for prompt C but improved performance for prompt B, confirming that noise effectiveness depends on prompt context rather than intrinsic seed quality.
While Experiment 1 shows how noise effectiveness varies across prompts, the following two experiments examine whether similar patterns also emerge within a fixed prompt.
Experiment 2: Intra-Prompt Attention Consistency.
Table 2: Pairwise distances within attention maps.
Type High (intra) Low (intra) Cross
Euclid. 1.635 1.567 1.735
To examine whether BANSA uncertainty relates to attention stability, we generate five samples per prompt (total 100 prompts), selected the highest- and lowest-scoring seeds, and computed pairwise Euclidean distances between their attention maps (Table 2). Low-BANSA seeds showed smaller intra-group distances, indicating more stable and consistent patterns, while high-BANSA seeds exhibited higher variability.
Experiment 3: Latent Trajectory and Expressiveness.
Table 3: Latent stability and visual dynamics.
Group Traj. Var. $\downarrow$ Var. $\uparrow$
High BANSA 52.34 0.041
Low BANSA 51.07 0.053
Beyond attention maps, we further investigate latent-space dynamics. Across 100 prompts with five repetitions each, we compare high- and low-BANSA seeds over all denoising steps (Table 3). We measured two metrics: (1) Latent Trajectory Variation, obtained by applying a Butterworth low-pass filter to isolate low-frequency structural components (Wu et al., 2024; Yuan et al., 2025) and computing their temporal variation, where lower values indicate smoother and more stable trajectories; and (2) Intra-frame Variance, the average spatial variance within frames, where higher values indicate richer dynamic expressiveness. Low-BANSA seeds showed both lower trajectory variation and higher intra-frame variance, combining temporal stability with more expressive video generation.
Takeaway and Relation to Prior Work.
The three findings indicate that BANSA identifies seeds with lower uncertainty, leading to (1) prompt-specific improvements, (2) more stable attention, and (3) smoother yet more expressive latent dynamics. This is consistent with FreeInit Wu et al. (2024) and FreeQPrior (Yuan et al., 2025), which emphasize stable latent initialization and low-frequency structure. BANSA complements these insights by offering an uncertainty-driven selector that explains why such seeds generalize within a prompt while remaining prompt-dependent across prompts.
@@ 5 Experiments
Experimental Setting. We evaluate ANSE on diverse T2V diffusion backbones—AnimateDiff (Guo et al., 2024b), CogVideoX-2B/5B (Yang et al., 2024), Wan2.1 (Team, 2025), and HunyuanVideo (Kong et al., 2024). To rigorously evaluate our method, we follow each model’s official sampling protocol; for resolution, Wan2.1 is evaluated at 480p and HunyuanVideo at 360p due to resource limits. Results for noise-prior refinement baselines (FreqPrior (Yuan et al., 2025)) are reported only on AnimateDiff, as they lack official support on the others and incur $\sim\!3\times$ inference time. ANSE is orthogonal and can be combined with these priors. Unless noted, we use a noise pool of M=10M{=}10 with K=10K{=}10 stochastic passes per seed, and Bernoulli-masked attention with p=0.2p{=}0.2. All experiments run on NVIDIA H100 GPUs. Further details are in the Appendix C.
Evaluation Metric. We evaluate with VBench (Huang et al., 2024), which reports a quality score and a semantic score, combined into a total score on a 0–100 scale. For AnimateDiff and CogVideoX-2B/5B, we report both quality and semantic scores. For HunyuanVideo and Wan2.1, we restrict evaluation to the six quality dimensions due to computational constraints, while noting that semantic evaluation can also be applied in principle. All reported results are obtained by repeating each evaluation dimension five times and averaging the scores.
Table 4: Quantitative results on VBench using AnimateDiff, CogVideoX-2B and -5B.
@@
Backbone Model Method Quality Score Semantic Score Total Score Inference Time
AnimateDiff (Guo et al., 2024b) Vanilla 80.22 69.03 77.98 28.23s
+ Ours 81.66 71.09 79.33 31.33s(+10.98%)
Freqprior 81.22 70.45 79.07 58.01(+105.36%)
+ Ours 82.23 73.23 80.43 61.12s(+5.36%)
CogVideoX-2B (Yang et al., 2024) Vanilla 82.08 76.83 81.03 247.8s
+ Ours 82.56 78.06 81.66 269.3s(+8.67%)
CogVideoX-5B (Yang et al., 2024) Vanilla 82.53 77.50 81.52 667.3s
+ Ours 82.70 78.10 81.71 754.1s(+13.1%)
Table 5: Quantitative results on quality score metrics for HunyuanVideo and Wan2.1
Backbone Model Method Subject Consistency $\uparrow$ Background Consistency $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Temporal Flickering $\uparrow$ Inference Time $\downarrow$
HunyuanVideo (Kong et al., 2024) Vanilla 0.9562 0.9656 0.9850 0.6276 0.6137 0.9921 52.3s
+ Ours 0.9612 0.9661 0.9858 0.6268 0.6151 0.9938 59.8s(+14.34%)
Wan2.1 (Team, 2025) Vanilla 0.9562 0.9656 0.9850 0.6276 0.6137 0.9943 43.8s
+ Ours 0.9612 0.9661 0.9858 0.6268 0.6151 0.9956 50.7s(+15.75%)
@@
Figure 4: Qualitative comparison with and without ANSE. (CogvideoX-2B, 5B and Wan2.1). With ANSE, videos exhibit improved visual quality, better text alignment, and smoother motion transitions compared to the baseline. White numbers show BANSA scores for the same prompt.
Quantitative Comparison. As shown in Table 4, ANSE consistently improves performance across diverse T2V backbones. On AnimateDiff, ANSE outperforms the vanilla baseline and further demonstrates clear advantages over noise-initialization methods such as FreqPrior (Yuan et al., 2025).
Notably, ANSE is also fully compatible with FreqPrior, achieving even higher scores when combined. Table 4 also reports results on CogVideoX-2B and -5B, which adopt the more advanced MMDiT architecture. ANSE improves quality, semantic, and total VBench scores on both scales, demonstrating robust generalization across architectures and model sizes.
Table 6: Comparison with FVMD metrics on MSRVTT dataset.
Backbone Model Methods FVMD $\downarrow$
HunyuanVideo (Kong et al., 2024) Vanilla 17641.19
+ Ours 16491.68
Wan2.1(1.3B) (Team, 2025) Vanilla 16495.59
+ Ours 14306.19
Table 5 presents additional results on recent state-of-the-art models, HunyuanVideo and Wan2.1. ANSE achieves consistent improvements across quality metrics, underscoring its plug-and-play applicability to a wide spectrum of video diffusion models—from U-Net to MMDiT, and from mid-scale to the latest state-of-the-art systems. The statistical signicanse analysis of these are provied in Appendix J.
We further evaluate motion quality using the Fréchet Video Motion Distance (FVMD) (Liu et al., 2024) on MSR-VTT (Xu et al., 2016), as shown in Table 6. We sample 200 text prompts from the test split and generate videos with ANSE, using the corresponding real videos as the reference distribution. ANSE consistently obtains lower FVMD scores, indicating improved motion fidelity and reinforcing the motion-related gains observed in VBench.
Qualitative Comparison. Figure 4 presents representative qualitative comparisons on CogVideoX-2B, CogVideoX-5B, and Wan2.1 with and without ANSE. Our approach consistently enhances semantic fidelity, motion realism, and visual clarity across diverse prompts. For example, in separate prompts such as "exploding" and "descends gracefully", ANSE captures critical semantic transitions—generating visible explosions in the former and preserving smooth temporal continuity in the latter. In other prompts like "koala playing the piano" and "tasting beer", it produces anatomically coherent bodies with natural, expressive motion. These examples highlight ANSE’s ability to improve spatial-temporal fidelity while generalizing effectively to large-scale video diffusion models. Additional qualitative results, including other backbones, are provided in Appendix L.
Computational Cost. As shown in Table 4 and Table 5, ANSE introduces only a minimal increase in inference time while delivering consistent quality gains. On AnimateDiff, inference time increases by just +10.98%, compared to more than +105% when combined with FreqPrior and over +200% for FreeInit. On CogVideoX models, the overhead is similarly small: +8% on the 2B variant and +13% on the 5B variant. For more recent architectures, ANSE adds only +14% on HunyuanVideo and +15% on Wan2.1. The overhead comes only from seed evaluation and leaves the sampling process and memory usage unchanged. In contrast, FreeInit and FreqPrior require multiple full passes, causing large slowdowns. ANSE improves quality across diverse backbones while keeping inference overhead below +15%, making it a practical plug-and-play solution for video diffusion.
6 Ablation Study
Comparison of Acquisition Functions. Using CogVideoX-2B, we compare BANSA with random sampling, entropy-based selection, and two variants: BANSA (B) with Bernoulli masking and BANSA (D) with dropout-based stochasticity (Table 8). All methods improve over the baseline, but BANSA (B) consistently achieves the best scores across quality, semantic, and total metrics. This suggests that Bernoulli masking better captures attention-level uncertainty than dropout, underscoring the importance of modeling uncertainty in line with the model’s structure.
Effect of Ensemble Size $K$. We investigate how the number of stochastic forward passes $K$ influences subject and background consistency (Table 8). Both metrics improve as $K$ increases from 1 to 10, indicating that larger ensembles provide more reliable noise evaluation. Performance saturates at $K=10$, which we set as the default in all experiments.
Effect of Noise Pool Size $M$. We analyze the effect of noise pool size $M$, which controls the diversity of candidate seeds in BANSA. Larger $M$ improves the chance of finding high-quality seeds but increases inference cost. As shown in Figure 8 (Appendix), performance saturates around $M=10$ for CogVideoX-2B and -5B, which we adopt as the default for each model.
@@ 6 Ablation Study
Table 10: Component analysis of Bernoulli masking parameters on Wan2.1 1.3B.
Backbone Model Bernoulli $p$ Subject Consistency $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Overall Consistency $\uparrow$
Wan2.1 1.3B 0.0 0.9254 0.9751 0.6464 0.6459 0.2731
0.2 0.9310 0.9786 0.6559 0.6516 0.2763
0.5 0.9308 0.9779 0.6543 0.6533 0.2727
0.7 0.9302 0.9778 0.6557 0.6529 0.2726
1.0 0.9232 0.9753 0.6451 0.6508 0.2721
Effect of Bernoulli masking probability $p$.
We conduct a sensitivity analysis on the Bernoulli masking probability $p$ using the Wanx 2.1 (1.3B) model. As shown in Table 10, both disabling masking ($p=0$) and fully random masking ($p=1$) degrade performance, indicating that the masking probability has a meaningful impact on the results. The method remains stable within the intermediate range $p\in[0.2,0.7]$, where overall performance varies only slightly. Among these configurations, $p=0.2$ delivers the best performance, which supports our choice of the default setting.
Table 11: Quantitative evaluation on the IPOC-2B after post-training and Wan2.1 14B model.
Backbone Model Method Subject Consistency $\uparrow$ Background Consistency $\uparrow$ Temporal Flickering $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Dynamic Degree $\uparrow$
IPOC-2B Vanilla 0.9273 0.9535 0.9918 0.9741 0.6554 0.664 0.7527
+ Ours 0.9281 0.9534 0.9921 0.9745 0.6561 0.6644 0.7721
Wan2.1 14B Vanilla 0.9290 0.9587 0.9931 0.9759 0.6572 0.6552 0.5916
+ Ours 0.9302 0.9596 0.9938 0.9772 0.6601 0.6529 0.6277
Evaluation on strong video backbones and post-RL models.
We further evaluate our method on stronger settings, including the large-capacity Wan2.1 14B backbone and the post-trained IPOC-2B model. IPOC-2B is refined with a DPO-related preference alignment method (Yang et al., 2025), making it a competitive post-RL baseline. As shown in Table 11, our method consistently improves over the vanilla models across most VBench dimensions, demonstrating generalization to larger architectures and robustness under preference-aligned post-RL training.
User Study.
We conducted a human preference study in which evaluators compared paired videos generated with and without our method along two criteria: overall quality and text–video alignment. As shown in Fig. 5 (a), our method is consistently preferred on both aspects, indicating improvements in perceptual quality and semantic alignment. The details are provided in Appendix K.
Analysis on correlation with actual quality.
We examined how BANSA scores relate to quality by sampling diverse prompts and noise inputs and visualizing their score–quality trends in Fig. 5(b). Pearson correlations show strong negative relationships with Subject Consistency (–0.6672), Background Consistency (–0.7492), and Motion Smoothness (–0.6769), indicating that lower BANSA scores reliably associate with higher-quality and more stable generations.
Figure 5: (a) User study showing consistent human preference for our method in overall quality and text–video alignment. (b) Correlation analysis where BANSA scores exhibit strong negative trends with key VBench metrics, indicating that lower scores correspond to higher-quality generations.
@@ 7 Discussion and Limitations
Our method focuses on noise–seed selection based on model uncertainty, which introduces several limitations. As shown in Figure 6, even seeds with low BANSA scores may still produce unnatural results because ANSE does not modify the subsequent sampling process. Moreover, BANSA captures uncertainty at the attention level, which does not fully account for semantic or aesthetic quality. Although evaluating multiple candidates with stronger quality metrics could reduce this gap, such an approach is computationally prohibitive. A promising direction for future work is to integrate ANSE with post-training refinement methods—such as Self-Forcing (SF) (Huang et al., 2025)—that directly enhance sample quality. Notably, ANSE already works effectively with post-RL refinement methods like IPoC, and extending this compatibility to SF represents a natural next step. Such combinations may yield complementary gains in both quality and robustness beyond seed selection alone.
@@ 8 Conclusion
We present ANSE, a framework for active noise selection in video diffusion models. At its core is BANSA, an acquisition function that leverages attention-derived uncertainty to identify noise seeds yielding confident, consistent attention and thus higher-quality generations. BANSA adapts the BALD principle to the generative setting by operating in attention space, with efficient deployment enabled through Bernoulli-masked attention and lightweight layer selection. Experiments across multiple T2V backbones show that ANSE improves video quality and prompt alignment with minimal inference overhead. This introduces an inference-time scaling paradigm, enhancing generation not by altering the model or sampling steps, but through informed seed selection.
Acknowledgement
This work was partially supported by Institute of Information &\& Communications Technology Planning &\& Evaluation (IITP) grant funded by the Korean government (MSIT) (Artificial Intelligence Graduate School Program (GIST)) (No. 2019-0-01842).
@@ Appendix A Supplementary Section
  • Proof of Proposition 1 from the main paper regarding the BANSA Zero condition in Section B.
  • Implementation details of ANSE and evaluation metrics in Section C.
  • Further explanation of layer selection through cumulative BANSA correlation in Section D.
  • Additional ablation studies, including full-layer BANSA score analysis (Section E), temporal scope effects (Section F), and attention-space effects (Section G).
  • Additional analysis, including effect of CFG scale (Section H), equivalent computation budget (Section I), statistical validation (Section J), and user study protocol (Section K).
  • Additional qualitative results demonstrating the impact of BANSA Score in Section L.
Appendix B Proof of Proposition 1